### v2

# The Energy Loss Phenomenon in RLHF: A New Perspective on Mitigating Reward Hacking

Abstract
This work identifies the Energy Loss Phenomenon in Reinforcement Learning from Human Feedback (RLHF) and its connection to reward hacking. Specifically, energy loss2 in the final layer of a Large Language Model (LLM) gradually increases during the RL process, with an excessive increase in energy loss characterizing reward hacking. Beyond empirical analysis, we further provide a theoretical foundation by proving that, under mild conditions, the increased energy loss reduces the upper bound of contextual relevance in LLMs, which is a critical aspect of reward hacking as the reduced contextual relevance typically indicates overfitting to reward model-favored patterns in RL. To address this issue, we propose an Energy loss-aware PPO algorithm (EPPO) which penalizes the increase in energy loss in the LLM’s final layer during reward calculation to prevent excessive energy loss, thereby mitigating reward hacking. We theoretically show that EPPO can be conceptually interpreted as an entropy-regularized RL algorithm, which provides deeper insights into its effectiveness. Extensive experiments across various LLMs and tasks demonstrate the commonality of the energy loss phenomenon, as well as the effectiveness of EPPO in mitigating reward hacking and improving RLHF performance.
Reward Hacking, Reward Overoptimization, Reinforcement Learning from Human Feedback, Large Language Models
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is a key technique for aligning Large Language Models (LLM) with human preferences, enabling helpful, honest, and harmless responses (Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022). This process involves Supervised Fine-Tuning (SFT) on human-annotated data, followed by Reinforcement Learning (RL) that maximizes rewards from a learned proxy Reward Model (RM) based on human preference rankings. Such approaches have greatly improved LLM alignment with human expectations, driving advanced AI applications (Touvron et al., 2023; Ouyang et al., 2022).
Figure 1: Illustration of the energy loss phenomenon. During response generation, the energy loss [^2] The energy loss of a given layer is defined as the difference between the $L_{1}$ norms of its input and output hidden states. in RLHF model’s final layer tends to be higher than that in SFT models; ❷ hacking RLHF responses exhibit excessive energy loss compared to normal ones.
Despite the success, RLHF faces a challenge known as reward hacking, where the policy model optimization under the proxy RM diverges from true human preference (Gao et al., 2023). The issue arises because the RM is an imperfect proxy for human preferences (Miao et al., 2024). LLMs can exploit these imperfections by generating well-formed but less relevant responses that overfit to reward model-favored patterns, such as excessive redundancy or caution, to achieve high RM scores through manipulation (Coste et al., 2024; Chen et al., 2024b). Such overfitting prioritizes adherence to reward model-favored patterns, rather than accurately capturing user intent, which often undermines the contextual relevance of responses (Zhang et al., 2024). Some typical hacking samples are provided in Appendix K.
Recent efforts have primarily focused on mitigating reward hacking by enhancing reward modeling (Coste et al., 2024; Chen et al., 2024b; Miao et al., 2024) or designing RL regularizations (Singhal et al., 2024; Ouyang et al., 2022). While reward modeling strategies show promise, challenges such as overfitting, misspecification, and misgeneralization make achieving accurate and robust reward modeling a formidable task in practice (Casper et al., 2023; Azar et al., 2024). These issues highlight the need for tailored regularization techniques to address reward hacking in RLHF effectively.
However, existing RL regularization techniques addressing reward hacking mainly focus on imposing constraints on the output space, such as Kullback-Leibler (KL) divergence and response length penalties (Singhal et al., 2024; Ouyang et al., 2022), overlooking the underlying mechanisms of reward hacking. As a result, they lack deep understanding of network behavior, inevitably restricting the optimization landscape of the policy model and often compromising RLHF performance (Azar et al., 2024; Chen et al., 2024b).
In this work, we aim to uncover the underlying mechanisms of reward hacking within LLMs333 Our work focuses on policy models. For clarity, the term “LLMs” in this paper refers specifically to policy models in RLHF. for developing more effective RL regularization techniques. To this end, we investigate the internal representation dynamics of LLMs during the RL process and closely examine the generation of samples prone to hacking. We empirically observe a consistent increase in energy loss in the LLM’s final layer during the RL process, with an excessive increase in energy loss characterizing reward hacking. We formally define the energy loss phenomenon in Definition 2 and provide a visual illustration in Figure 1. This observation aligns with recent advancements in LLM representation engineering, highlighting the critical role of the LLM’s final layer in encapsulating essential information (Zhang et al., 2024).
Besides empirical analysis, we further establish a theoretical foundation for the energy loss phenomenon. Given the complexity of reward hacking, our analysis focuses on a critical aspect of reward hacking—the contextual relevance of responses. In Theorem 3, we prove that under mild conditions, the energy loss in the LLM’s final layer is negatively correlated with the upper bound of contextual mutual information, thus increased energy loss suppresses the contextual relevance of responses. This trend typically indicates overfitting to reward model-favored patterns and serves as a critical aspect of reward hacking. To validate Theorem 3, we empirically show in Section 6.2 that this suppression is consistently observed across different LLMs.
We demonstrate the widespread presence of the energy loss phenomenon on four popular LLMs (Llama3-8B (Dubey et al., 2024), Llama2-7B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), and Deepspeek-7B (Bi et al., 2024)) and two representative tasks (general dialogue (Bai et al., 2022) and summarization (Stiennon et al., 2020)). To address this phenomenon, we propose an Energy loss-aware PPO algorithm (EPPO) that mitigates reward hacking by penalizing the increase in energy loss in the LLM’s final layer during reward calculation. Unlike existing RLHF algorithms that mitigate reward hacking by imposing direct constraints on output spaces, EPPO takes a distinct approach by concentrating solely on the energy loss in the final layer. This design enables a broader scope for policy exploration and optimization, leading to superior performance, as shown in Figure 2. Additionally, in Section 6.1, we theoretically show that EPPO can be conceptually interpreted as an entropy-regularized RL algorithm, providing deeper insights into its effectiveness. Our main contributions are as follows:
Figure 2: Response comparison between EPPO and baselines on AlpacaFarm dataset, as assessed by GPT-4. By focusing solely on energy loss in the LLM’s final layer, EPPO benefits from a broader optimization landscape, outperforming existing RLHF algorithms that directly constrain output spaces, such as PPO with length penalty (PPO w/ LP) and PPO with KL penalty (PPO w/ KL).
             LLM: Llama3-8B         LLM: Mistral-7B            LLM: DeepSeek-7B
Refer to caption Refer to caption Refer to caption
Figure 3: Energy loss in the final layer of various LLMs during PPO and EPPO training, including Llama3-8B, Mistral-7B, and DeepSeek-7B. Observations: Energy loss in the LLM’s final layer gradually increases as RL progresses. By penalizing the increase in energy loss , EPPO effectively suppresses the excessive energy loss increase during RL process.
$\bullet$ We empirically identify the energy loss phenomenon, where reward hacking internally manifests in LLMs as an excessive increase in energy loss in the final layer.
$\bullet$ We theoretically support our findings by proving that, under mild conditions, increased energy loss suppresses contextual relevance in LLMs—a key aspect of reward hacking.
$\bullet$ We propose EPPO, an energy loss-aware PPO algorithm that mitigates reward hacking by penalizing the increase in energy loss in LLM’s final layer during reward calculation.
$\bullet$ We demonstrate the effectiveness of EPPO in mitigating reward hacking and improving RLHF performance across four popular LLMs and two representative tasks.
2 Preliminaries
In this section, we introduce the formulation of RLHF and discuss the reward hacking phenomenon in RLHF.
2.1 Reinforcement Learning from Human Feedback
Recently, RLHF has emerged as a standard technique for aligning LLM behavior with human preferences (Ouyang et al., 2022; Bai et al., 2022). When fine-tuning an LLM with RL, a KL penalty is often added to prevent significant deviations from the initial policy. Denoting $x$ and $y$ as the prompt from the dataset $\mathcal{D}$ and the corresponding response, the optimization objective is given by (Ouyang et al., 2022):
$\arg\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\left[r(y|x)-\beta\log\left(\frac{\pi_{\theta}(y|x)}{\pi^{\text{SFT}}(y|x)}\right)\right],$
where $\pi_{\theta}(\cdot)$ is the policy model, $\pi^{\text{SFT}}(\cdot)$ is the SFT model, $r(\cdot)$ is the RM, and $\beta$ is the KL penalty weight. In this work, we adopt the industry-standard RL algorithm, Proximal Policy Optimization (PPO) (Schulman et al., 2017), to optimize the policy $\pi_{\theta}(\cdot)$ under the given objective (Ouyang et al., 2022; Touvron et al., 2023; Bai et al., 2022).
2.2 Reward Hacking in RLHF
In RLHF, an RM is employed to approximate human preferences, eliciting the need for human feedback during RL. However, since the RM serves as a proxy for human preferences, its optimization does not always align with genuine human goals. In practice, RLHF optimization continuously improves proxy reward performance. However, alignment with true human preferences is typically achieved only in the early training stage, after which RLHF performance often deteriorates—a phenomenon known as reward hacking.
Given the challenges of achieving robust reward modeling in real-world scenarios (Casper et al., 2023; Wang et al., 2024), designing RL regularizations to mitigate reward hacking is crucial (Ouyang et al., 2022). Existing regularizations often target output space, such as KL divergence and response length penalties (Singhal et al., 2024; Ouyang et al., 2022), failing to address the internal mechanisms of reward hacking in LLMs. This inevitably limits the optimization landscape and often degrades RLHF performance (Azar et al., 2024; Chen et al., 2024b). In this paper, we delve into the underlying mechanisms of reward hacking and develop more effective RL regularizations for reward hacking mitigation.
         LLM: Llama3-8B & RLHF Algorithm: PPO          LLM: Mistral-7B & RLHF Algorithm: PPO          LLM: DeepSeek-7B & RLHF Algorithm: PPO
Refer to caption Refer to caption Refer to caption
         LLM: Llama3-8B & RLHF Algorithm: EPPO          LLM: Mistral-7B & RLHF Algorithm: EPPO          LLM: DeepSeek-7B & RLHF Algorithm: EPPO
Refer to caption Refer to caption Refer to caption
Figure 4: Energy loss distribution of SFT and RLHF model responses on the AlpacaFarm dataset. Hacking samples are identified by GPT-4, following the protocol in Miao et al. (2024), with further details provided in Section 5.1. Rows correspond to RLHF algorithms (PPO and EPPO), and columns to LLMs (Llama3-8B, Mistral-7B, and DeepSeek-7B). Observations: Hacking samples from RLHF models exhibit a more excessive increase in energy loss compared to normal ones, as shown in the top row. EPPO effectively mitigates reward hacking by suppressing the increase in energy loss during RL, as evidenced by the disappearance of the red bins in the bottom row.
3 The Energy Loss Phenomenon
In this section, we introduce the energy loss phenomenon, an underlying mechanism of reward hacking, supported by empirical observations of energy loss dynamics in LLMs during RL and their connection to reward hacking. Furthermore, we provide a theoretical foundation for our findings.
Definition 1.
Given an input $x$, let $h_{\ell}^{\text{in}}(x)$ and $h_{\ell}^{\text{out}}(x)$ represent the input and output hidden states of the $\ell$-th layer of the LLM, respectively. The energy loss in the $\ell$-th layer during response generation is defined as:
$\Delta E_{\ell}(x)=\|h_{\ell}^{\text{in}}(x)\|_{1}-\|h_{\ell}^{\text{out}}(x)\|_{1},$
where the energy of a hidden state is measured by $L_{1}$-norm.
Definition 1 quantifies the energy loss in LLM layers during a single forward pass. For text generation tasks, it is calculated as the average energy loss across all tokens in the response, as each token requires a separate forward pass.
Building on our empirical observations that energy loss increases during RL training, with excessive increase indicating reward hacking, as detailed in Section 3.1, we formally define the energy loss phenomenon as:
Definition 2.
An algorithm exhibits the energy loss phenomenon if: ❶ energy loss in the LLM’s final layer gradually increases during the RL process, and ❷ an excessive increase in energy loss indicates reward hacking.
3.1 Empirical Evidence on Energy Loss Phenomenon
In this section, we present empirical evidence that supports the two primary characteristics of the energy loss phenomenon. Specifically, our analysis focuses on three representative LLMs: Llama3-8B, Mistral-7B, and DeepSeek-7B, as well as a widely-used dataset, AlpacaFarm. Additional models, such as Llama2-7B, and datasets, including Anthropic-Helpful, Anthropic-Harmless, and Reddit TL;DR, are provided in Appendixes B and C.
We first examine the energy loss in the LLM’s final layer during RL training, as illustrated in Figure 3. We can observe that the energy loss in the LLM’s final layer gradually increases during the RL process. This pattern is consistent across various LLMs and datasets, as shown in Appendix B, underscoring the widespread presence of the energy loss phenomenon ❶ in Definition 2.
In Figure 4, we illustrate the relationship between energy loss and reward hacking. Hacking samples are identified by GPT-4 following Miao et al. (2024), detailed in Section 5.1. As observed, while RLHF models generally exhibit higher energy loss than SFT models, hacking samples from RLHF models show a more excessive increase in energy loss compared to normal ones. This consistent pattern across various LLMs and datasets, as shown in Appendix C, shows the commonality of energy loss phenomenon ❷ in Definition 2.
Table 1: Response comparison on Llama3-8B under GPT-4 evaluation between EPPO and existing strategies addressing reward hacking, categorized into RL and RM algorithms (denoted as Cate.), highlighting EPPO’s effectiveness in enhancing RLHF performance.
Cate. Opponent Anthropic-Helpful Anthropic-Harmless AlpacaFarm TL;DR Summary
  Win     /     Tie     /     Lose   Win     /     Tie     /     Lose   Win     /     Tie     /     Lose   Win     /     Tie     /     Lose
RL SFT 72%19%9% 69%20%11% 62%24%14% 81%10%9%
PPO 76%17%7% 78%12%10% 68%20%12% 66%21%13%
PPO w/ KL 59%30%11% 46%33%21% 51%29%20% 57%25%18%
PPO w/ LP 69%22%9% 52%30%18% 58%26%16% 46%32%22%
RM ERM-Mean 64%27%9% 66%23%11% 57%27%16% 52%28%20%
ERM-WCO 48%37%15% 63%24%13% 45%32%23% 56%26%18%
ERM-UWO 59%31%10% 64%24%12% 49%30%21% 40%36%24%
WARM 56%33%11% 57%28%15% 53%28%19% 43%34%23%
Table 2: Response comparison on Llama3-8B under GPT-4 evaluation among EPPO, advanced reward modeling techniques, and their combinations, demonstrating EPPO’s compatibility with reward modeling techniques to further enhance RLHF performance.
Comparison Anthropic-Helpful Anthropic-Harmless AlpacaFarm TL;DR Summary
  Win     /     Tie     /     Lose   Win     /     Tie     /     Lose   Win     /     Tie     /     Lose   Win     /     Tie     /     Lose
EPPO vs. ODIN 61%28%11% 44%35%22% 56%28%16% 35%39%26%
EPPO+ODIN vs. ODIN 67%25%8% 52%30%18% 62%26%12% 42%36%22%
EPPO vs. InfoRM 40%39%21% 48%32%20% 39%36%25% 38%37%25%
EPPO+InfoRM vs. InfoRM 46%37%17% 60%25%15% 44%34%22% 44%36%20%
3.2 Theoretical Analysis on Energy Loss Phenomenon
We establish a theoretical foundation for the energy loss phenomenon by examining the response-context relevance—a critical aspect of reward hacking (Zhang et al., 2024).
Theorem 3.
Let $X$ and $Y$ be the input and output random variables of an LLM. For any layer $\ell$, let $H^{\text{in}}_{\ell}$ and $H^{\text{out}}_{\ell}$ be its input and output hidden state random variables. Define the global energy offset as $\alpha_{\ell}=\mathbb{E}[\|H^{\text{out}}_{\ell}\|_{1}]-\mathbb{E}[\|H^{\text{in}}_{\ell}\|_{1}]$. The contextual mutual information $I(X;Y)$ satisfies:
$I(X;Y)\leq\frac{1}{2\sigma^{2}}\mathbb{E}\left[\left(\Delta E_{\ell}(X)+\alpha_{\ell}\right)^{2}\right]+\log\sqrt{2\pi\sigma^{2}},$
where $\sigma$ is a positive constant.
Theorem 3 bounds $I(X;Y)$ using energy loss $\Delta E_{\ell}(x)$ and offset $\alpha_{\ell}$, showing energy loss’s effect on transmitting contextual information. The proof is provided in Appendix A.
Building on Theorem 3 and utilizing the properties of quadratic functions, we derive the following corollary:
Corollary 4.
For an arbitrary layer $\ell$ in the LLM, when the energy loss $\Delta E_{\ell}(x)$ falls below the global energy offset $-\alpha_{\ell}$, it exhibits a negative correlation with the upper bound of the contextual mutual information $I(X;Y)$.
Corollary 4 examines scenarios where the increase in energy loss remains below a specific threshold, revealing that such an increase reduces the upper bound of contextual mutual information and thereby suppresses contextual relevance—a critical aspect in reward hacking. However, our theory suggests that when the energy loss surpasses this threshold, its impact on contextual relevance remains an open question, as the increased energy loss instead raises the upper bound.
Remark I: In Section 6.2, we empirically demonstrate the negative correlation between energy loss in the LLM’s final layer and the contextual relevance of responses, validating our theoretical analysis and the energy loss phenomenon.
Table 3: Response comparison between EPPO and RL algorithms addressing reward hacking on different LLMs under GPT-4 evaluation, demonstrating the consistent advantage of EPPO in enhancing RLHF performance across diverse LLMs..
LLM Opponent Anthropic-Helpful Anthropic-Harmless AlpacaFarm TL;DR Summary
  Win     /     Tie     /     Lose   Win     /     Tie     /     Lose   Win     /     Tie     /     Lose   Win     /     Tie     /     Lose
Llama2 SFT 57%25%18% 65%19%16% 47%32%21% 78%12%10%
PPO 65%22%13% 72%17%11% 54%28%18% 65%24%11%
PPO w/ KL 49%29%22% 56%23%21% 41%33%26% 55%28%17%
PPO w/ LP 54%26%20% 39%36%25% 49%31%20% 51%30%19%
Mistral SFT 61%22%17% 56%27%17% 56%32%12% 76%11%13%
PPO 48%29%23% 62%23%15% 50%35%15% 67%16%17%
PPO w/ KL 38%36%26% 36%30%34% 41%38%21% 49%30%21%
PPO w/ LP 45%32%23% 47%27%26% 45%36%19% 46%32%22%
DeepSeek SFT 76%15%9% 75%13%12% 59%29%12% 72%17%11%
PPO 72%18%10% 78%12%10% 56%31%13% 64%23%13%
PPO w/ KL 59%29%12% 46%31%23% 47%36%17% 51%34%15%
PPO w/ LP 66%24%10% 54%28%18% 52%33%15% 47%36%17%
4 Energy Loss-Aware PPO (EPPO)
Our analysis in Section 3 highlights the energy loss phenomenon in RLHF, revealing that reward hacking manifests internally as excessive energy loss in the LLM’s final layer. Building on this observation, we propose the Energy loss-aware PPO (EPPO) algorithm to mitigate reward hacking.
The core idea behind EPPO is to penalize the energy loss increase in reward calculations, aiming to suppress excessive energy loss increase and mitigate reward hacking. Given a prompt $x$ sampled from the dataset $\mathcal{D}$, the energy loss in the SFT model’s final layer $\Delta E^{\text{SFT}}_{final}(x)$ is precomputed and stored. During the RL process, the corresponding energy loss in RLHF model’s final layer $\Delta E^{\text{RLHF}}_{final}(x)$ is computed. The energy loss increase/variation [^4] During RL, energy loss variation is empirically equivalent to its increase, as we have demonstrated that energy loss consistently increases throughout the RL process in Figure 3. is calculated as $|\Delta E^{\text{SFT}}_{final}(x)-\Delta E^{\text{RLHF}}_{final}(x)|$ and then used to penalize the reward from the reward model to suppress the excessive energy loss increase. With this energy loss-based penalty, the optimization objective of EPPO is formulated as:
$\arg\max_{\pi_{\theta}}\mathbb{E}_{x\sim\mathcal{D},y\sim\pi_{\theta}(\cdot|x)}\left[\hat{r}(y|x)\right],$ (1)
$\hat{r}(y|x)$ $=r(y|x)-\eta\left|\Delta E^{\text{SFT}}_{final}(x)-\Delta E^{\text{RLHF}}_{final}(x)\right|,$
where $\eta$ is the trade-off parameter, with its sensitivity analysis in Appendix G and experimental setups in Appendix J.1. Figure 3 illustrates the energy loss dynamics during the RL using PPO and EPPO. By penalizing the increase in energy loss, EPPO significantly suppresses the excessive energy loss increase. Furthermore, Figure 4 demonstrates the effectiveness of EPPO in mitigating reward hacking.
Remark II: Our EPPO can be conceptually interpreted as an entropy-regularized RL algorithm (Ahmed et al., 2019; Haarnoja et al., 2017), offering benefits like improved stability and enhanced exploration, as discussed in Section 6.1.
5 Experiments
In this section, we highlight EPPO’s effectiveness in mitigating reward hacking and enhancing RLHF performance across various LLMs and tasks.
5.1 Setup
Model and Training Data. In our experiments, we evaluate EPPO on four popular LLMs, including Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B, Llama2-7B (Touvron et al., 2023), Mistral-7B (Jiang et al., 2023), and Deepspeek-7B (Bi et al., 2024), and two typical tasks, i.e., general dialogue task and the summarization task. For general dialogue, base models are fine-tuned on the ShareGPT dataset [^6] https://huggingface.co/datasets/anon8231489123/ShareGPT_Vicuna_unfiltered following prior works (Miao et al., 2024; Zheng et al., 2024; Rafailov et al., 2024). Reward modeling is performed using the Anthropic-HH dataset (Bai et al., 2022), and instructions from this dataset are also used for RL optimization. For summarization, we use the Reddit TL;DR dataset (Stiennon et al., 2020) for SFT, reward modeling, and RL. More details are provided in Appendix J.1.
Baseline. In addition to the Supervised Fine-Tuning model (SFT) and the RLHF model using the PPO algorithm (PPO), the baselines in our experiments also include two PPO algorithms proposed for mitigating reward hacking: PPO with KL penalty (PPO w/ KL(Ouyang et al., 2022) and PPO with length penalty (PPO w/ LP(Singhal et al., 2024). Additionally, six reward modeling methods addressing reward hacking are included: ERM-Mean (Coste et al., 2024), ERM-UWO (Coste et al., 2024), ERM-WCO (Coste et al., 2024), WARM (Rame et al., 2024), ODIN (Chen et al., 2024b), and InfoRM (Miao et al., 2024). Unless otherwise specified, all reward modeling methods are optimized with the PPO algorithm during RL. Details and hyper-parameter setting of the baselines are provided in Appendix J.2.
Evaluation Data. For general dialogue task, we use in-distribution data (test set of Anthropic-HH (Bai et al., 2022), including Anthropic-Helpful and Anthropic-Harmless datasets) and out-of-distribution data (test set of AlpacaFarm (Dubois et al., 2024)). For summarization task, the Reddit TL;DR test set (Stiennon et al., 2020) is used.
GPT-4 Evaluation. We evaluate EPPO by comparing the win ratio of RLHF model responses under EPPO against baselines. GPT-4 is adopted for assessment due to its strong alignment with human evaluations (Chen et al., 2023; Zheng et al., 2024). To mitigate positional bias (Wang et al., 2018), each sample pair is evaluated twice with reversed order. We use the GPT-4 prompt with the highest human agreement in AlpacaEval (Li et al., 2023), as detailed in Appendix J.3. To ensure the reliability of our experiments, we also adopt Claude-3.5-Sonnet and human as evaluators in Appendix F.
GPT-4 Identification of Hacking Samples. To explore the relationship between energy loss and reward hacking, we use AI feedback to identify hacking samples following Miao et al. (2024). Specifically, we first outline common hacking behaviors (Coste et al., 2024; Zhai et al., 2023) and then use GPT-4 to evaluate responses based on these criteria, with prompts detailed in Appendix J.3 and over 95% human agreement validated in Appendix I.
         LLM: Llama3-8B          LLM: Mistral-7B          LLM: DeepSeek-7B
Refer to caption Refer to caption Refer to caption
Figure 5: Win rate dynamics of RLHF models compared to SFT models during RL process across various LLMs on AlpacaFarm dataset under GPT-4 evaluation. Win rate is calculated as $win+0.5*tie$ for more accurate performance assessment. Observations: Comparison methods either show limited RL performance gains or significantly degrade in later RL stages indicating reward hacking. In contrast, EPPO effectively mitigates reward hacking while significantly boosting final RLHF performance.
Refer to caption Refer to caption
RLHF Algorithm: PPO RLHF Algorithm: EPPO
Figure 6: T-SNE visualization of response distributions within InfoRM’s latent space on AlpacaFarm and Llama3-8B, where hacking responses are identified as outliers (Miao et al., 2024). Observation: EPPO markedly reduces outliers in InfoRM’s latent space, showcasing its effectiveness in mitigating reward hacking.
5.2 Main Results of RLHF Performance
Tables 1 and 2 compare the RLHF performance of EPPO with other baselines. Key findings include: EPPO outperforms existing RLHF algorithms proposed for mitigating reward hacking. We conjecture that this is because EPPO focuses solely on the final layer’s energy loss, thereby alleviating the compromises to the policy optimization landscape induced by existing output space-constrained RL algorithms, i.e., PPO w/ KL [^7] PPO w/ KL under different KL penalties are compared in Appendix H to ensure fairness and reliability of our experiments. and PPO w/ LP. EPPO surpasses reward modeling approaches addressing reward hacking. While existing methods significantly improve RLHF performance by enhancing the robustness of reward modeling, they may still be vulnerable to spurious features in reward modeling and distribution shifts in RL stage, potentially causing a certain degree of reward hacking (Casper et al., 2023). A detailed discussion of these baselines is provided in Appendix E. EPPO integrates seamlessly with reward modeling approaches to provide complementary benefits. Compared with recent reward modeling strategies like ODIN and InfoRM, our EPPO exhibits better RLHF performance and achieves further improvement when combined with these methods, as shown in Table 2.
To validate the effectiveness of EPPO across different models, we extend our evaluation beyond Llama3-8B. In Table 3, we further compare the RLHF performance of EPPO with existing RLHF algorithms on three additional LLMs: Llama2-7B, Mistral-7B, and DeepSeek-7B. The results demonstrate that EPPO consistently achieves enhanced RLHF performance across various LLMs.
5.3 Main Results of Reward Hacking Mitigation
Given the unavailability of the gold score, we demonstrate the effectiveness of EPPO in mitigating reward hacking from the following three perspectives:
$\bullet$ GPT-4 Identification. As detailed in Section 5.1, we identify hacking samples by summarizing common hacking behaviors (Miao et al., 2024) and using AI feedback (i.e., GPT-4) to assess whether an RLHF response qualifies as hacking; please refer to Appendix J.3 for prompt details. Figure 4 illustrates the distribution of GPT-4-judged normal and hacking responses from RLHF models trained with PPO and EPPO. The results demonstrate that by penalizing energy loss, EPPO effectively prevents reward hacking. Additional results validating EPPO’s effectiveness across various LLMs and datasets are provided in Appendix D.1.
$\bullet$ GPT-4 Win Rate Dynamics in RL. Following Rafailov et al. (2024), we further assess reward hacking mitigation by tracking GPT-4 win rate dynamics during RL. Results across different LLMs are provided in Figure 5. As observed, PPO performance deteriorates significantly in later RL stages, indicating reward hacking. By constraining response length, PPO w/ LP improves RL training stability, but still suffers from reward hacking to some extent, as evidenced by a decline in the later stages. This is because response length fails to capture all patterns like excessive caution, exampled in Appendix L. While PPO w/ KL effectively mitigates reward hacking by constraining KL divergence in the output space, it inevitably restricts policy exploration, thus limiting RLHF performance gains. In contrast, by constraining only the energy loss in the LLM’s final layer, EPPO not only effectively mitigates reward hacking but also enables a broader policy optimization landscape, leading to a notable boost in final RLHF performance. Additional results on more datasets are provided in Appendix D.2.
$\bullet$ Representation Analysis using InfoRM. Beyond the above two GPT-4-based evaluations, we also utilize the InfoRM technique (Miao et al., 2024), which identifies hacking responses as significant outliers in its latent space. Figure 6 visualizes the response distribution from SFT and RLHF models on Llama3-8B using the AlpacaFarm dataset. The results demonstrate that EPPO significantly reduces outliers in InfoRM’s latent space, effectively preventing hacking responses from the RLHF model. Additional visualizations on more datasets are provided in Appendix D.3.
         LLM: Llama3-8B          LLM: Mistral-7B          LLM: DeepSeek-7B
Refer to caption Refer to caption Refer to caption
Figure 7: The relationship between energy loss in the LLM’s final layer and response-context relevance, measured by the contextual dependency strength metric proposed for general dialogue task (Chen et al., 2024a). Observation: Excessively increased energy loss suppresses the contextual relevance of responses, which aligns with our theoretical analysis that significant energy loss increases tighten the upper bound of contextual relevance, potentially diminishing it in practice.
6 Discussion
6.1 EPPO vs. Entropy-Regularized RL
This section explores the connection between EPPO and entropy-regularized RL algorithms (Ahmed et al., 2019; Haarnoja et al., 2017). In Corollary 5, we theoretically establish the relationship between energy loss and output entropy. The detailed proof is provided in Appendix A.
Corollary 5.
Under the setting of Theorem 3, the energy loss $\Delta E_{\ell}(X)$ also contributes to an upper bound for the output entropy $\mathcal{H}(Y)$, expressed as:
$\mathcal{H}(Y)\leq\frac{1}{2\sigma^{2}}\mathbb{E}\left[\left(\Delta E_{\ell}(X)+\alpha_{\ell}\right)^{2}\right]+\log\sqrt{2\pi\sigma^{2}}.$
Thus, if the energy loss stays below a specific threshold, it negatively correlates with the upper bound of output entropy.
According to Corollary 5, by suppressing excessive energy loss increase in the LLM’s final layer, our EPPO also prevents LLM’s output entropy from dropping excessively. From this perspective, EPPO can be conceptually interpreted as an entropy-regularized RL algorithm, inherently offering advantages such as improved stability and enhanced exploration capability (Ahmed et al., 2019; Haarnoja et al., 2017). This interpretation provides a theoretical basis for the superior performance of EPPO.
6.2 Energy Loss vs. Response-Context Relevance
In Section 3.2, we theoretically demonstrate that when energy loss remains below a specific threshold, it exhibits a negative correlation with the upper bound of the contextual relevance of model responses, forming the theoretical foundation for the identified energy loss phenomenon. To validate this, we empirically investigate the relationship between energy loss in the LLM’s final layer and the contextual relevance of its responses, which is quantified using the contextual dependency strength (CDS) metric (Chen et al., 2024a). CDS, specifically designed for dialogue tasks, is defined as the perplexity difference of a response with and without context. More details are provided in Appendix J.4.
Figure 7 illustrates the relationship between energy loss and response-context relevance for three representative LLMs. The results reveal that excessively increased energy loss indeed suppresses response-context relevance, aligning with our theoretical analysis: a substantial increase in energy loss significantly tightens the upper bound of contextual relevance, potentially reducing contextual relevance in practice. This suggests that the LLM may overfit reward model-favored patterns, leading to reward hacking. These findings support the validity of the energy loss phenomenon.
7 Related Work
7.1 Reward Hacking Mitigation in RLHF
Existing approaches to mitigating reward hacking in RLHF fall into two main categories: enhancing reward model (RM) robustness and designing regularization techniques to constrain policy updates. The first includes scaling RMs (Gao et al., 2023), using RM ensembles (Coste et al., 2024; Eisenstein et al., 2024), composing RMs from multiple perspectives (Moskovitz et al., 2024; Rame et al., 2024), optimizing RM datasets (Zhu et al., 2024a), applying information bottleneck theory (Miao et al., 2024), and addressing length bias (Shen et al., 2023; Chen et al., 2024b). The second focuses on adding regularization terms like KL divergence or response length during optimization (Singhal et al., 2024; Touvron et al., 2023; Ouyang et al., 2022). These complementary strategies target different aspects of reward hacking and together enhance final RLHF performance.
In this paper, we aim to understand the underlying mechanism of reward hacking by analyzing LLM’s internal states and propose tailored regularization techniques for mitigation. Unlike existing methods that constrain the output space (e.g., KL or response length penalties), our approach focuses solely on regularizing the energy loss in the LLM’s final layer. This enables a broader policy optimization space and achieves more effective reward hacking mitigation.
7.2 Representation Engineering for Neural Networks
Neural networks, often perceived as chaotic and opaque, have been shown to develop semantically meaningful internal representations. Early work on word embeddings reveals semantic relationships, compositionality (Mikolov et al., 2013), and biases within text corpora (Bolukbasi et al., 2016). Subsequent studies demonstrate that text embeddings could encode dimensions like commonsense morality without explicit instruction (Schramowski et al., 2019). Recent research has increasingly focused on regulating LLM behavior through internal representations to address attributes such as honesty, fairness, and harmlessness (Zou et al., 2023; Park et al., 2024; Azaria & Mitchell, 2023; Ji et al., 2024; Zhu et al., 2024b). Building on this foundation, our paper shows that reward hacking in RLHF manifests within LLM representations and introduces a tailored mitigating strategy.
8 Conclusion
This study identifies the energy loss phenomenon as an underlying mechanism of reward hacking, characterized by a gradual increase in LLM’s final-layer energy loss during RL, where an excessive energy loss empirically corresponds to reward hacking. We provide a theoretical foundation, proving that increased energy loss suppresses contextual response relevance, a critical aspect of reward hacking. To address this, we propose EPPO, an energy loss-aware PPO algorithm that incorporates an energy loss-based penalty into rewards. Unlike RLHF methods that constrain LLM output space, EPPO focuses on final-layer energy loss, enabling a broader scope for policy exploration and optimization. We also theoretically show that EPPO can be interpreted as an entropy-regularized RL algorithm, offering theoretical insights into its effectiveness. Extensive experiments confirm the widespread occurrence of the energy loss phenomenon and underscore the superior performance of EPPO.
Impact Statement
In reinforcement learning from human feedback, reward hacking refers to the phenomenon where the optimization of a policy model, while seemingly effective under a proxy reward model, deviates from true human objectives. This divergence often reduces the helpfulness of large language models in various ways, ranging from generating less meaningful content to exhibiting excessive caution in responses. In this work, we identify the energy loss phenomenon in RLHF and its connection to reward hacking. Building on this discovery, we propose an energy loss-aware PPO algorithm that significantly mitigates reward hacking and enhances RLHF performance. Our research is dedicated to better aligning large models with human preferences, thereby increasing their positive contributions to human society. Consequently, this study raises no ethical concerns and does not pose any adverse effects on society.
Appendix A Proof of Theorem and Corollary
Theorem 3. Let $X$ and $Y$ be the input and output random variables of an LLM. For any layer $\ell$, let $H^{\text{in}}_{\ell}$ and $H^{\text{out}}_{\ell}$ be its input and output hidden state random variables. Define the global energy offset as $\alpha_{\ell}=\mathbb{E}[\|H^{\text{out}}_{\ell}\|_{1}]-\mathbb{E}[\|H^{\text{in}}_{\ell}\|_{1}]$. The contextual mutual information $I(X;Y)$ satisfies:
$I(X;Y)\leq\frac{1}{2\sigma^{2}}\mathbb{E}\left[\left(\Delta E_{\ell}(X)+\alpha_{\ell}\right)^{2}\right]+\log\sqrt{2\pi\sigma^{2}},$
where $\sigma$ is a positive constant.
Proof: As defined in Definition 1, given an input $x$ to the LLM, the input and output hidden states of the $\ell$-th layer are denoted as $h_{\ell}^{in}(x)$ and $h_{\ell}^{out}(x)$, respectively. For simplicity, in the following proof, we abbreviate them as $h_{\ell}^{in}$ and $h_{\ell}^{out}$, representing instances of the random variables $H_{\ell}^{in}$ and $H_{\ell}^{out}$, respectively. Additionally, the mapping of the $\ell$-th layer of the LLM is denoted as $f_{\ell}(\cdot)$.
Since $f_{\ell}(\cdot)$ is a deterministic mapping, given an input hidden state $h_{\ell}^{in}$, the conditional distribution of the output hidden state $h_{\ell}^{out}$ is a Dirac distribution:
$p(h_{\ell}^{out}\mid h_{\ell}^{in})=\delta(h_{\ell}^{out}-f_{\ell}(h_{\ell}^{in})),$ (2)
where $\delta(\cdot)$ denotes the Dirac delta function, satisfying:
[…]
By introducing a variational distribution $q(h_{\ell}^{out})$, a variational upper bound for $I(H_{\ell}^{in};H_{\ell}^{out})$ can be derived as follows:
$I(H_{\ell}^{in};H_{\ell}^{out})$ $=\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log\frac{p(h_{\ell}^{in},h_{\ell}^{out})}{p(h_{\ell}^{in})p(h_{\ell}^{out})}\right]$ (4)
$=\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log\frac{\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))}{p(h_{\ell}^{out})}\right]$
$=\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log\frac{\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))q(h_{\ell}^{out})}{q(h_{\ell}^{out})p(h_{\ell}^{out})}\right]$
$=\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log\frac{\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))}{q(h_{\ell}^{out})}\right]-KL(p(h_{\ell}^{out})\|q(h_{\ell}^{out}))$
$\leq\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log\frac{\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))}{q(h_{\ell}^{out})}\right]$
$=\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))-\log q(h_{\ell}^{out})\right]$
The inequality in the above equation is derived from the non-negativity of the KL divergence. Based on the inequality $\log x\leq x-1$, we have:
$I(H_{\ell}^{in};H_{\ell}^{out})$ $\leq\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))-\log q(h_{\ell}^{out})\right]$ (5)
$\leq\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))-1-\log q(h_{\ell}^{out})\right]$
$=\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))\right]-1-\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log q(h_{\ell}^{out})\right]$
Next, we focus on simplifying $\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\delta(h_{\ell}^{out}-f(h_{\ell}^{in}))\right]$:
[…]
Substituting Equation (6) into Equation (5), we obtain:
$I(H_{\ell}^{in};H_{\ell}^{out})$ $\leq-\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log q(h_{\ell}^{out})\right]$ (7)
To make the variational upper bound in Equation (7) as tight as possible, we need $q(h_{\ell}^{out})$ to closely approximate $p(h_{\ell}^{out})$ (Poole et al., 2019). Let $\alpha_{\ell}=\mathbb{E}[\|H^{\text{out}}_{\ell}\|_{1}]-\mathbb{E}[\|H^{\text{in}}_{\ell}\|_{1}]$, and we make the following assumption about the variational distribution $q(h_{\ell}^{out})$:
$q(h_{\ell}^{out})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{1}{2\sigma^{2}}(\|h_{\ell}^{out}\|_{1}-\mathbb{E}[\|H_{\ell}^{in}\|_{1}]-\alpha_{\ell})^{2}\right),$ (8)
where $\sigma$ is a positive constant.
The variational upper bound of the mutual information $I(H_{\ell}^{in};H_{\ell}^{out})$ is now expressed as:
$I(H_{\ell}^{in};H_{\ell}^{out})$ $\leq-\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\log q(h_{\ell}^{out})\right]$ (9)
$=\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\frac{1}{2\sigma^{2}}(\|h_{\ell}^{out}\|_{1}-\mathbb{E}[\|H_{\ell}^{in}\|_{1}]-\alpha_{\ell})^{2}\right]+\log\sqrt{2\pi\sigma^{2}}$
Since $g(x)=(x-k)^{2}$ is a convex function, the variational upper bound in Equation (9) can be further refined using Jensen’s inequality as:
$I(H_{\ell}^{in};H_{\ell}^{out})$ $\leq\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\frac{1}{2\sigma^{2}}(\|h_{\ell}^{out}\|_{1}-\mathbb{E}[\|H_{\ell}^{in}\|_{1}]-\alpha_{\ell})^{2}\right]+\log\sqrt{2\pi\sigma^{2}}$ (10)
$\leq\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\frac{1}{2\sigma^{2}}(\|h_{\ell}^{out}\|_{1}-\|h_{\ell}^{in}\|_{1}-\alpha_{\ell})^{2}\right]+\log\sqrt{2\pi\sigma^{2}}$
$=\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\frac{1}{2\sigma^{2}}(\|h_{\ell}^{in}\|_{1}-\|h_{\ell}^{out}\|_{1}+\alpha_{\ell})^{2}\right]+\log\sqrt{2\pi\sigma^{2}}$
Let $X$ and $Y$ denote the input and output of the LLM, respectively. Since $X\rightarrow H_{\ell}^{in}\rightarrow H_{\ell}^{out}\rightarrow Y$ forms a Markov chain, the data processing inequality gives:
$I(X;Y)\leq I(H_{\ell}^{in};H_{\ell}^{out})\leq\mathbb{E}_{p(h_{\ell}^{in},h_{\ell}^{out})}\left[\frac{1}{2\sigma^{2}}(\|h_{\ell}^{in}\|_{1}-\|h_{\ell}^{out}\|_{1}+\alpha_{\ell})^{2}\right]+\log\sqrt{2\pi\sigma^{2}}$ (11)
Combining with Definition 1, we derive the final version of the variational upper bound for $I(X;Y)$:
$I(X;Y)\leq\mathbb{E}\left[\frac{1}{2\sigma^{2}}(\Delta E_{\ell}(X)+\alpha_{\ell})^{2}\right]+\log\sqrt{2\pi\sigma^{2}}.$ (12)
Thus, Theorem 3 holds.
  Corollary 5. Under the setting of Theorem 3, the energy loss $\Delta E_{\ell}(X)$ also contributes to an upper bound for the output entropy $\mathcal{H}(Y)$, expressed as:
$\mathcal{H}(Y)\leq\frac{1}{2\sigma^{2}}\mathbb{E}\left[\left(\Delta E_{\ell}(X)+\alpha_{\ell}\right)^{2}\right]+\log\sqrt{2\pi\sigma^{2}}.$
Thus, when the energy loss stays below a specific threshold, it is negatively correlated with the output entropy.
Proof: Since $X$ and $Y$ represent the input and output of the LLM, respectively, $Y$ is uniquely determined by $X$, which implies $\mathcal{H}(Y|X)=0$. Now, we have:
$I(X;Y)=\mathcal{H}(Y)-\mathcal{H}(Y|X)=\mathcal{H}(Y)$ (13)
By Equation (13) and Theorem 3, Corollary 5 holds.
Appendix B More Results on Energy Loss Dynamics.
In this section, we further examine the trends in energy loss in the LLM’s final layer during the RL process across additional models and tasks, corresponding to the first characteristic of the energy loss phenomenon defined in Definition 2. Results from four widely-used LLMs and two representative tasks are presented in Figure 8. The key findings are as follows: ❶ Across all models and tasks, the energy loss in the LLM’s final layer consistently demonstrates a gradual upward trend throughout the RL process, underscoring the widespread presence of the first characteristic of the energy loss phenomenon as defined in Definition 2. ❷ In all settings, our EPPO effectively suppresses the excessive growth in energy loss, which is consistent with both our motivation for developing EPPO.
                      LLM: Llama3-8B   &  Task: General Dialogue                       LLM: Llama3-8B   &  Task: Summerization
       Refer to caption        Refer to caption
                      LLM: Llama2-7B   &  Task: General Dialogue                       LLM: Llama2-7B   &  Task: Summerization
       Refer to caption        Refer to caption
                      LLM: Mistral-7B   &  Task: General Dialogue                       LLM: Mistral-7B   &  Task: Summerization
       Refer to caption        Refer to caption
                      LLM: DeepSeek-7B   &  Task: General Dialogue                       LLM: DeepSeek-7B   &  Task: Summerization
       Refer to caption        Refer to caption
Figure 8: Energy loss in the final layer of various LLMs during the RL process using PPO and our EPPO algorithms. From top to bottom: The LLMs shown are Llama3-8B, Llama2-7B, Mistral-7B, and DeepSeek-7B, respectively. From top to bottom: The tasks are general dialogue and Summerization tasks, respectively.
Appendix C More Results on Energy Loss Distribution.
In this section, we further analyze the energy loss distribution of responses before and after RLHF, as well as its relationship to reward hacking, across a diverse range of LLMs and datasets. This analysis corresponds to the second characteristic of the energy loss phenomenon defined in Definition 2; see Figures 9, 10, 11, and 12 for related results. The key findings are: ❶ The excessive increase in energy loss consistently corresponds to reward hacking across all LLMs and datasets, emphasizing the widespread presence of the second characteristic of the energy loss phenomenon in Definition 2.Our EPPO effectively mitigates reward hacking in all settings by suppressing the excessive increase in energy loss, aligning with our motivation for developing EPPO and the hacking mitigation analysis presented in Section 5.3. Notably, despite our best efforts, hacking samples in the summarization task are challenging to identify with simple natural language instruction for GPT-4 precisely. Therefore, unlike the general dialogue task, where GPT-4 serves as the hacking annotator, we rely on the recently proposed representation-based detection method, InfoRM, to identify hacking samples in the summarization task. Related human evaluation results are detailed in Section I.
C.1 Energy Loss Distribution on Llama3-8B
                   Dataset: AlpacaFarm   &   RLHF Algorithm: PPO                    Dataset: AlpacaFarm   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Anth.-Helpful   &   RLHF Algorithm: PPO                    Dataset: Anth.-Helpful   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Anth.-Harmless   &   RLHF Algorithm: PPO                    Dataset: Anth.-Harmless   &    RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Reddit TL;DR   &   RLHF Algorithm: PPO                    Dataset: Reddit TL;DR   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
Figure 9: Energy loss distribution of responses from the SFT model and RLHF model on Llama3-8B. The RLHF model responses are categorized as normal or hacking responses, as judged by GPT-4 for general dialogue task and by InfoRM for summarization task. From left to right: The RLHF algorithms utilized are PPO and EPPO, respectively. From top to bottom: The datasets used for response generation are AlpacaFarm, Anthropic-Helpful, Anthropic-Harmless, and Reddit TL;DR datasets, respectively.
C.2 Energy Loss Distribution on Mistral-7B
                   Dataset: AlpacaFarm   &   RLHF Algorithm: PPO                    Dataset: AlpacaFarm   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Anth.-Helpful   &   RLHF Algorithm: PPO                    Dataset: Anth.-Helpful   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Anth.-Harmless   &   RLHF Algorithm: PPO                    Dataset: Anth.-Harmless   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Reddit TL;DR   &   RLHF Algorithm: PPO                    Dataset: Reddit TL;DR   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
Figure 10: Energy loss distribution of responses from the SFT model and RLHF model on Mistral-7B. The RLHF model responses are categorized as normal or hacking responses, as judged by GPT-4 for general dialogue task and by InfoRM for summarization task. From left to right: The RLHF algorithms utilized are PPO and EPPO, respectively. From top to bottom: The datasets used for response generation are AlpacaFarm, Anthropic-Helpful, Anthropic-Harmless, and Reddit TL;DR datasets, respectively.
C.3 Energy Loss Distribution on Llama2-7B
                   Dataset: AlpacaFarm   &   RLHF Algorithm: PPO                    Dataset: AlpacaFarm   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Anth.-Helpful   &   RLHF Algorithm: PPO                    Dataset: Anth.-Helpful   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Anth.-Harmless   &   RLHF Algorithm: PPO                    Dataset: Anth.-Harmless   &    RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Reddit TL;DR   &   RLHF Algorithm: PPO                    Dataset: Reddit TL;DR   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
Figure 11: Energy loss distribution of responses from the SFT model and RLHF model on Llama2-7B. The RLHF model responses are categorized as normal or hacking responses, as judged by GPT-4 for general dialogue task and by InfoRM for summarization task. From left to right: The RLHF algorithms utilized are PPO and EPPO, respectively. From top to bottom: The datasets used for response generation are AlpacaFarm, Anthropic-Helpful, Anthropic-Harmless, and Reddit TL;DR datasets, respectively.
C.4 Energy Loss Distribution on DeepSeek-7B
                   Dataset: AlpacaFarm   &   RLHF Algorithm: PPO                    Dataset: AlpacaFarm   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Anth.-Helpful   &   RLHF Algorithm: PPO                    Dataset: Anth.-Helpful   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Anth.-Harmless   &   RLHF Algorithm: PPO                    Dataset: Anth.-Harmless   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
                   Dataset: Reddit TL;DR   &   RLHF Algorithm: PPO                    Dataset: Reddit TL;DR   &   RLHF Algorithm: EPPO
       Refer to caption        Refer to caption
Figure 12: Energy loss distribution of responses from the SFT model and RLHF model on DeepSeek-7B. The RLHF model responses are categorized as normal or hacking responses, as judged by GPT-4 for general dialogue task and by InfoRM for summarization task. From left to right: The RLHF algorithms utilized are PPO and EPPO, respectively. From top to bottom: The datasets used for response generation are AlpacaFarm, Anthropic-Helpful, Anthropic-Harmless, and Reddit TL;DR datasets, respectively.
Appendix D More Results on Reward Hacking Mitigation of EPPO
In this section, we further validate the effectiveness of our EPPO in mitigating reward hacking across additional LLMs and datasets, focusing on the following three aspects: GPT-4 identification, GPT-4 win rates dynamics in RL, and representation analysis using InfoRM.
D.1 GPT-4 Identification
We have presented the distribution of normal and hacking responses from RLHF models trained with PPO and our EPPO across a wide range of LLMs and datasets in Figures 9, 10, 11, and 12. As observed, our EPPO consistently demonstrates significant effectiveness in mitigating reward hacking across all settings, consistent with our analysis in Section 5.3
D.2 GPT-4 Win Rates Dynamics in RL
In Section 5.3, we presented the GPT-4 win rate dynamics on the AlpacaFarm dataset during the RL process across various LLMs. In this section, we further illustrate the GPT-4 win rate dynamics during RL on additional evaluation datasets, using Llama3-8B as an example. Related results are presented in Figure 13. As observed, PPO exhibits a sharp decline in GPT-4 win rate during the later stages of training, indicating the occurrence of reward hacking. By constraining the KL divergence and limiting response length in the output space of the LLM, PPO w/ KL and PPO w/ LP significantly improve the stability of RL training but at the cost of reduced policy exploration, leading to limited RLHF performance gains. In contrast, by solely constraining the energy loss in the LLM’s final layer, our EPPO enables a broader policy optimization landscape, effectively mitigating reward hacking while achieving superior final RLHF performance. These findings are consistent with the conclusions drawn in Section 5.3, further validating the effectiveness of our approach.
         Dataset: AlpacaFarm          Dataset: Anthropic Helpful
         Refer to caption          Refer to caption
         Dataset: Anthropic Harmless          Dataset: Reddit TL;DR
         Refer to caption          Refer to caption
Figure 13: The win rate dynamics of RLHF model compared to SFT model during RL training on Llama3-8B, under GPT-4 evaluation. To better measure performance, we calculate the win rate as $win+0.5*tie$. From left to right and from top from bottom: The evaluation datasets are AlpacaFarm, Anthropic Helpful, Anthropic Harmless, and Reddit TL;DR datasets, respectively.
D.3 Representation Analysis using InfoRM
In this section, we further compare the response distributions of PPO and our EPPO in the latent space of InfoRM across additional evaluation datasets. According to Miao et al. (2024), outliers in the latent space of InfoRM typically correspond to hacking samples. The related visualizations are presented in Figure 14. As observed, compared to PPO, our EPPO consistently mitigates the occurrence of outliers across all datasets, effectively overcoming reward hacking. This observation aligns with our proposed solution discussed in Section 5.3.
            Dataset: AlpacaFarm & RLHF Algorithm: PPO             Dataset: AlpacaFarm & RLHF Algorithm: EPPO
            Refer to caption             Refer to caption
            Dataset: Anth. Helpful & RLHF Algorithm: PPO             Dataset: Anth. Helpful & RLHF Algorithm: EPPO
            Refer to caption             Refer to caption
            Dataset: Anth. Harmless & RLHF Algorithm: PPO             Dataset: Anth. Harmless & RLHF Algorithm: EPPO
            Refer to caption             Refer to caption
            Dataset: Reddit TL;DR & RLHF Algorithm: PPO             Dataset: Reddit TL;DR & RLHF Algorithm: EPPO
            Refer to caption             Refer to caption
Figure 14: T-SNE visualization of the response distribution in the latent space of InfoRM from the SFT and RLHF models on the Llama3-8B. According to Miao et al. (2024), outliers in the latent space of InfoRM typically correspond to hacking samples. From left to right: The RLHF algorithms utilized are PPO and EPPO, respectively. From top to bottom: The datasets used for response generation are AlpacaFarm, Anthropic-Helpful, Anthropic-Harmless, Reddit TL;DR datasets, respectively.
Appendix E Reward Hacking Mitigation of More Compared Methods
In Appendix D.2, we have analyzed reward hacking mitigation in constrained RLHF algorithms using GPT-4 win rate dynamics. Here, we extend our analysis to reward modeling methods, including ERM-Mean, ERM-WCO, ERM-UWO, and WARM. Due to budget constraints, we employ the recently proposed representation-based reward hacking detection technique, InfoRM, which aligns closely with GPT-4 evaluations (Miao et al., 2024). According to Miao et al. (2024), hacking samples often appear as significant outliers in InfoRM’s latent space after RLHF. Figure 15 shows the response distribution in InfoRM’s latent space. While these reward modeling methods significantly reduce hacking samples by improving reward model robustness, they still show some vulnerability. In contrast, by constraining energy loss in the LLM’s final layer, our method (Standard RM + EPPO) more effectively mitigates reward hacking.
RM: Standard RM & RLHF Algorithm: PPO RM: ERM-Mean & RLHF Algorithm: PPO RM: ERM-WCO & RLHF Algorithm: PPO
Refer to caption Refer to caption Refer to caption
RM: Standard RM & RLHF Algorithm: EPPO RM: ERM-UWO & RLHF Algorithm: PPO RM: WARM & RLHF Algorithm: PPO
Refer to caption Refer to caption Refer to caption
Figure 15: T-SNE visualization of the response distribution in the latent space of InfoRM from the SFT and RLHF models on Llama3-8B and the AlpacaFarm dataset. Observations: While reward modeling methods reduce hacking samples compared to Standard RM, they still show some vulnerability, whereas our method (Standard RM + EPPO) more effectively mitigates reward hacking.
Appendix F More Win-Tie-Lose Results under Various Evaluators
To ensure the reliability of our experiments, in addition to GPT-4 evaluations, we provide results from Claude-3.5 (Sonnet) and human evaluations in Table 4, using Llama3-8B and representative datasets for general dialogue and summarization tasks as examples. For human evaluation, we presented responses generated by our method and the baselines to two expert annotators proficient in LLM alignment studies and fluent in English. The sources of these responses were anonymized to prevent bias. Human evaluators were asked to assess which response was more useful, harmless, and of higher quality. In cases of disagreement, the annotators reassessed their evaluations to reach a consensus. As shown in Table 4, our method consistently outperforms constrained RLHF algorithms proposed to address reward hacking across all evaluators.
Table 4: Comparison of win, tie, and loss ratios between our EPPO and existing constrained RLHF algorithms addressing reward hacking, evaluated on Llama3-8B under various evaluators, highlights the consistent advantages of EPPO in improving RLHF performance.
Evaluator Opponent AlpacaFarm TL;DR Summary
  Win      /      Tie      /      Lose   Win      /      Tie      /      Lose
GPT-4 PPO w/ KL 51%29%20% 57%25%18%
PPO w/ LP 58%26%16% 46%32%22%
Claude-3.5 PPO w/ KL 50%27%23% 55%21%24%
PPO w/ LP 56%23%21% 49%25%26%
Human PPO w/ KL 49%32%19% 51%29%20%
PPO w/ LP 54%29%17% 45%30%25%
Appendix G Sensitivity Analysis of Hyperparameter in EPPO
In this section, we evaluate the impact of the hyperparameter $\eta$ (the trade-off coefficient in Equation 1) on the RLHF performance of our EPPO, using Llama3-8B as a representative model. The corresponding results are shown in Figure 16. As observed, our EPPO is relatively robust to variations in the hyperparameter $\eta$ to some extent and achieves optimal performance on the general dialogue task and summary task when $\eta$ is set to 35 and 25, respectively.
Task: General Dialogue & Dataset: AlpacaFarm Task: Summarization & Dataset: Reddit TL:DR
Refer to caption Refer to caption
Figure 16: Win rate (%) of RLHF model compared to SFT model using our EPPO with different hyperparameter $\eta$, evaluated under GPT-4. To better measure performance, we calculate the win rate as $win+0.5*tie$.
Appendix H Comparison Results with Different KL Penalties
To ensure the fairness and reliability of the experiments, we also report the comparison results between our EPPO and PPO w/ KL under different KL penalties in Table 5. As shown, our EPPO consistently exhibits significant advantages, regardless of the KL penalty configurations. This phenomenon can be attributed to our method’s distinctive approach, which constrains only the energy loss in the LLM’s final layer. In contrast, PPO w/ KL directly regulates the KL divergence in the output space. As a result, our method facilitates a broader and more flexible policy optimization landscape.
Table 5: Comparison of win, tie, and lose ratios between our EPPO and PPO w/ KL with varying KL penalties, evaluated on Llama3-8B under GPT-4, demonstrates the consistent advantages of our EPPO in enhancing RLHF performance.
Method Opponent AlpacaFarm TL;DR Summary
  Win      /      Tie      /      Lose   Win      /      Tie      /      Lose
EPPO PPO w/ KL (kl=0.0001) 65%19%16% 64%23%13%
PPO w/ KL (kl=0.001) 56%27%17% 63%22%15%
PPO w/ KL (kl=0.01) 51%29%20% 61%23%16%
PPO w/ KL (kl=0.1) 58%24%18% 57%25%18%
PPO w/ KL (kl=0.5) 60%23%17% 63%23%14%
Appendix I Human Evaluation in Reward Hacking Identification
As mentioned in Section C, unlike the general dialogue task, where GPT-4 serves as the hacking annotator, we employ the representation-based InfoRM method for the summarization task, as hacking phenomena in summarization are challenging to accurately identify with simple natural language instructions for GPT-4.
In this section, we conduct a human evaluation to validate GPT-4 and InfoRM (Miao et al., 2024) as hacking annotators for the general dialogue and summarization tasks, respectively. Specifically, we randomly sampled 200 cases each from the AlpacaFarm dataset and the Reddit TL;DR dataset, which represent the general dialogue and summarization tasks, respectively. Two expert annotators, proficient in LLM alignment studies and fluent in English, evaluated these cases for hacking phenomena based on our predefined descriptions. In instances of disagreement, the annotators reassessed their evaluations to reach a consensus. For the general dialogue task, hacking phenomena were primarily characterized by deviations from user intent, excessive redundancy, and overly cautious responses. For the summarization task, hacking was identified as cases where the generated summaries lacked conciseness, included irrelevant information, or merely repeated the original text.
These human annotations served as references to calculate the accuracy of the GPT-4-based and InfoRM-based evaluators in identifying reward hacking. The results show a remarkable 97.5% agreement rate between human and GPT-4 evaluations for the general dialogue task and a 95% agreement rate between human and InfoRM evaluations for the summarization task, demonstrating the reliability of GPT-4-based hacking identification for general dialogue tasks and InfoRM-based identification for summarization tasks.
Appendix J Experiments Details
In this section, we provide more experimental details in this work.
J.1 Training Setup
Our experimental settings largely follow those outlined in Miao et al. (2024). SFT models were initialized from their respective pre-trained checkpoints. RMs are built upon SFT models, with the final layer removed and replaced by an additional linear layer to generate reward scores.
For fine-tuning during simulation experiments, training was conducted on a single node equipped with 8 A100-SXM80GB GPUs. We employed Data Parallelism (DP) and leveraged Automatic Mixed Precision (AMP) with bfloat16, utilizing the Deepspeed Zero framework (Rajbhandari et al., 2020). The training used a learning rate of 2e-5, with a single epoch for the SFT phase and a global batch size of 64.
In the reward modeling stage, a learning rate of 1e-6 was applied with a global batch size of 64, and models were trained on human preference datasets for a single epoch to minimize overfitting.
For the PPO training stage, the policy model was trained with a learning rate of 5e-7, while the critic model used a learning rate of 1e-6. Both were trained for a single epoch with a global batch size of 64. Sampling configurations included a temperature of 0.8, top-p of 0.9, and a maximum output token length of 512. The critic model was initialized from the SFT model weights, following recommendations from (Zheng et al., 2023). The Generalized Advantage Estimation parameter $\lambda$ was set to 0.95. Policy and critic optimizations were constrained by a clipping value of 0.2. The trade-off parameter in Equation (1) is selected from {$5i$, $i=1...8\}$ across all LLMs and tasks, manually adjusting to achieve optimal results.
J.2 Baseline Setup
J.2.1 RL Algorithms for Mitigating Reward Hacking
Proximal policy optimization (PPO) (Schulman et al., 2017). PPO is the core algorithm employed to achieve alignment with human preferences. In general dialogue and summarization tasks, we employ the reward model to train a policy separately that generates higher-quality responses.
PPO with Kullback-Leibler divergence penalty (PPO w/ KL) (Ouyang et al., 2022). To ensure that the output of the RLHF model does not significantly deviate from the distribution where the reward model remains accurate, researchers proposed imposing a KL divergence constraint in the output space of the LLM. In our experiments, the KL penalty coefficients were selected from {0.0001, 0.001, 0.01, 0.1, 0.5}, manually adjusting to achieve optimal results.
PPO with Length Penalty (PPO w/ LP) (Singhal et al., 2024). To mitigate the length bias, a specific manifestation of reward hacking, researchers proposed directly constraining response length. The response length penalty during reward calculation is formulated as: $\hat{r}(y|x)=r(y|x)+\left(1-\frac{\text{len}(y)}{N}\right)\sigma$, where $N$ represents the maximum allowable length, and $\sigma$ is a moving average of the batch reward standard deviation. In our experiments, $N$ is set to 250 and 100 for the general dialogue and summarization tasks, respectively, following the authors’ recommendations and our practical experience.
J.2.2 Reward Modeling Methods for Mitigating Reward Hacking
Ensemble RMs (ERM-Mean, ERM-WCO, and ERM-UWO) (Coste et al., 2024). To enhance the robustness of reward modeling, researchers proposed ensemble-based conservative optimization objectives, including mean optimization (Mean), worst-case optimization (WCO), and uncertainty-weighted optimization (UWO). In our experiments, five RMs are utilized, and the trade-off parameter $\lambda$ in UWO is selected from the candidates $\{0.5,0.1,0.01\}$, following the authors’ recommendations.
Weight Averaged RMs (WARM) (Rame et al., 2024). To further enhance the efficiency, reliability, and robustness of ensemble-based methods, researchers propose averaging multiple RMs. In our experiments, six RMs are averaged, following the authors’ recommendations.
Reward Disentangling Modeling (ODIN) (Chen et al., 2024b). Specifically addressing length bias in RLHF, researchers propose decoupling response length preference during RM training by employing two orthogonal linear heads. The head modeling response length preference is discarded during the inference stage. Following the authors’ recommendations, the hyperparameters $\lambda_{L}$ and $\lambda_{O}$ are both set to 1, in our experiments.
Infomation-Theoretic Reward Modeling (InfoRM) (Miao et al., 2024). To address the reward misgeneralization issue, researchers propose introducing a variational information bottleneck objective to filter out irrelevant information during RM training. In our experiments, following the authors’ recommendations, the latent space dimensionality is set to 128, and the trade-off parameter is set to 0.1.
J.3 GPT-4 Evaluation and Identification
In this section, we present the GPT-4 prompts used to compute win rate in general dialogue and summarization tasks, as well as the discriminator of hacking phenomenon. Detailed instructions provided to GPT-4 are illustrated in Figure 17.
Refer to caption Refer to caption Refer to caption
Figure 17: GPT-4 prompts used for general dialogue evaluation, summarization evaluation, and hacking sample identification.
J.4 Contextual Dependency Strength
In this section, we detail the calculation process of the contextual dependency strength metric, tailored for general dialogue scenarios and proposed in (Chen et al., 2024a), as follows:
Dependency $\text{Dependency Strength}(x,y)=\frac{\text{PPL}(y)-\text{PPL}(y|x)}{\text{PPL}(y)},$ (14)
where $x$ is the instruction and $y$ is the LLM response.
Appendix K Hacking Examples
To provide readers with an intuitive understanding of the hacking phenomenon, this section presents several representative hacking examples from general dialogue and summarization tasks, as shown in Figure 18.
In the general dialogue task, typical hacking behaviors include: ❶ The model refusing to answer a harmless user query due to excessive caution, instead generating irrelevant safety-related content, as demonstrated in the first sample of Figure 18. ❷ The model producing overly verbose responses, containing a significant amount of information unrelated to the user’s query, as illustrated in the third, fourth, and fifth samples of Figure 18.
In the summarization task, hacking phenomena are primarily characterized by overly detailed summaries that include unnecessary details or even repeat the original text, significantly deviating from the user’s intent for concise summaries, as shown in the last sample of Figure 18.
In general, a common characteristic of most hacking samples is weakened contextual relevance, where the model’s response gradually deviates from the user’s intent. This can be attributed to the model overfitting reward model-favored patterns, such as a preference for detailed or overly cautious responses, often at the expense of considering the preceding context and user intent.
Figure 18: Typical hacking samples in general dialogue and summarization tasks, with specific hacking segments highlighted in orange. The first four samples are from the general dialogue task, while the last is from summarization. A common characteristic of most hacking samples is weakened contextual relevance, i.e., the LLM’s response gradually deviates from the user’s intent.
Appendix L Qualitative Examples
In this section, we present some qualitative examples comparison in Figures 19, 20, and 21. We observe that due to constraints directly applied to the LLM’s output space, such as KL divergence and response length, PPO w/ KL and PPO w/ LP suffer from limited policy exploration space, resulting in restricted performance gains in RLHF. Moreover, as PPO w/ LP addresses only length bias, it remains susceptible to reward hacking (e.g., overly cautious responses) to some extent. In contrast, by solely constraining the energy loss in the LLM’s final layer, our EPPO effectively mitigates reward hacking while enabling a broader policy optimization space, leading to superior RLHF performance.
Figure 19: Qualitative Example I of RLHF model outputs with different RLHF algorithms. Specific hacking segments are highlighted in orange, while information covered by our EPPO but missed by competing methods is highlighted in green.
Figure 20: Qualitative example II of RLHF models output with different RLHF algorithms. Specific hacking segments are highlighted in orange, while information covered by our EPPO but missed by competing methods is highlighted in green.
Figure 21: Qualitative example III of RLHF models output with different RLHF algorithms. Specific hacking segments are highlighted in orange, while information covered by our EPPO but missed by competing methods is highlighted in green.